OpenStudy (anonymous):

need a bit of help setting this up. Find the volume of the ellipsoid x^2+y^2+9z^2=64. i know the volume is gonna be 4/3 pi r^3 and r here is 8

OpenStudy (anonymous):

Wow....I am so bad at this. sorry!

OpenStudy (anonymous):

that's ok, @PaulaLovesSchool13, i appreciate you taking the time to look and try. :D

OpenStudy (zpupster):

On the x-axis: For this case, y = z = 0, so we have x^2 = 64 and x = +or- 8. The distance from the center of the ellipsoid to either endpoint on the x axis is 8 On the y-axis: For this case, x = z = 0 y^2 = 64 so y = +or- 8 On the z-axis: For this case, x = y = 0 z^2= 64/9 z=8/3 so V=(4/3pi)(8)(8)(8/3)

OpenStudy (anonymous):

so we don't have to integrate for this? it's just as easy as that?

OpenStudy (ikram002p):

double integrate

ganeshie8 (ganeshie8):

familiar with change of vairables and jacobians ?

OpenStudy (anonymous):

jacobi what now?

OpenStudy (anonymous):

i know this is from the polar integrals section. never heard of that word

ganeshie8 (ganeshie8):

x^2+y^2+9z^2=64 (x/8)^2 + (y/8)^2 + (3z/8)^2 = 1

ganeshie8 (ganeshie8):

Oh you want to work it using a double integral is it ?

OpenStudy (anonymous):

well, zpupster had the right answer and that made sense to me, but i guess we're supposed to do it integrating. not sure why....more complicated that way.

ganeshie8 (ganeshie8):

zpupster has just used the volume of ellipsoid formula : \(\large \frac{4}{3}abc\)

ganeshie8 (ganeshie8):

\(\large \frac{4\pi}{3}abc\)

ganeshie8 (ganeshie8):

is this question from calc or geometry ?

OpenStudy (anonymous):

calc 3. polar integrals

ganeshie8 (ganeshie8):

double integrals / triple integrals ?

OpenStudy (anonymous):

either or. we've touched on triple.

ganeshie8 (ganeshie8):

good, then try setting up a triple integral

ganeshie8 (ganeshie8):

have they covered Jacobians yet ?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

brb.

OpenStudy (anonymous):

.....

ganeshie8 (ganeshie8):

it would be complicated without change of variables hmm

OpenStudy (zpupster):

wish i could help more i can understand the idea behind it but i can not explain it here are a couple of articles that may help you in understanding the triple integration as a solution for Volume of ellipsoid here is a tripe integral solution https://answers.yahoo.com/question/index?qid=20091010165412AArnjUe and here is a proof of the formula http://www.math24.net/calculation-of-volumes-using-triple-integrals.html

OpenStudy (anonymous):

ok. i'm just gonna move on to the next problem and talk to the teacher today. thanks for the help, guys.